# Sample-Efficient Online Learning in LM Agents via Hindsight Trajectory Rewriting

### v1

@@ 1 Introduction
While language models (LMs) have demonstrated remarkable generalization across tasks, their performance often degrades in unfamiliar or interactive environments, especially when learning from limited experience [Ziems et al., 2024, Kwa et al., 2025, Liang et al., 2023]. In such settings, sample efficiency becomes critical, particularly when interactions are costly (e.g., with humans or physical systems). For example, a conversational assistant deployed for the first time in a new organization likely does not know where to look for specific pieces of information, or the best means of communicating with specific people. Thus, creating agents that can learn and adapt to their environments over time is of critical importance in improving their everyday usability.
Here we study the problem of building efficient mechanisms for online learning in LM agents. We consider the setting where an LM agent receives queries one at a time in a streaming fashion. Existing LM agent frameworks typically approach this setting through reflection [Shinn et al., 2023, Zhao et al., 2024], memory [Wang et al., 2025b], or experience replay mechanisms [Zheng et al., 2024], which allow agents to revisit past episodes and improve over time. However, these methods primarily focus on storing or synthesizing experiences, and thus fail to fully exploit the LM’s ability to reason about counterfactuals—what could have led to success in past failures. This gap suggests an opportunity to design LM agents that actively rewrite and optimize their past experiences, converting failures into synthetic successes that improve future decision-making.
In this work, we introduce ECHO (Experience Consolidation via Hindsight Optimization), a framework that adapts hindsight experience replay (HER) to LM agents, enabling them to generate and learn from counterfactual trajectories for more sample-efficient learning. Our approach builds on hindsight experience replay methods from the RL literature [Andrychowicz et al., 2017, HER]. HER learns a goal-conditioned policy; during training, each attempt to reach a goal state $s$ that fails in an end state $s^{\prime}$ is interpreted as a successful trajectory for reaching $s^{\prime}$. For example, a trajectory in which an LM fails to slice an apple by attempting to grab a knife, dropping it, then grasping an apple may still be interpreted as a demonstration of a successful grasp. But HER and related methods are relatively limited in the set of trajectory modifications they can make—relabeling trajectories with goals, but not altering the structure demonstrated trajectories themselves.
@@
We test ECHO and various state-of-the-art agent architectures in stateful variants of a 2D GridWorld task (XMiniGrid, Nikulin et al. [2023]) and a question-answering task with multiple agents and tool calling (PeopleJoinQA, Jhamtani et al. [2025]). These environments require exploration to successfully solve all queries of the task. In XMiniGrid, the agent must explore to find various objects in different rooms, and in PeopleJoin, the agent has imprecise or incomplete information about which teammates have the information it needs to answer a question. We make these environments stateful by allowing agents to persist insights via a scratchpad memory. Next, we reset the environment to the same initial position or state and vary the queries or tasks posed to the agent. The agent can then infer information about the environment over time and improve its performance and efficiency.
On XMiniGrid-Stateful, ECHO outperforms the baseline reason-then-act (ReAct) LM agent [Yao et al., 2023] by 80% in average reward and the next best baseline by 42%. On PeopleJoinQA-Stateful, ECHO still outperforms the standard ReAct agent on both accuracy and efficiency, while being slightly worse than the best baseline in accuracy by 4.6% and tied in efficiency. We conclude that ECHO is a promising technique for improving the sample efficiency of language agents, especially in environments where rewards are sparse and the baseline language agent performs poorly.
@@ 2 Related Work
Language model (LM) agents are systems that use large language models as reasoning engines to interact with environments, make decisions, and execute actions over time [Yao et al., 2023, Schick et al., 2023]. These agents typically operate through a perception–action loop, where they observe their environment, reason about the current state, and generate actions [Wang et al., 2023]. Recent work has demonstrated LM agents’ capabilities across diverse domains, from web navigation and tool use to multi-agent collaboration and code generation [Liu et al., 2024, Qian et al., 2023, Jhamtani et al., 2025]. However, a key limitation of current LM agents is their reliance on static knowledge encoded during pre-training, making them less effective when deployed in novel environments that require exploration and adaptation [Zhou et al., 2024]. This motivates the need for agents that can accumulate experience and improve their performance through interaction with their environment over time.
@@
Following Sumers et al. [2024], we categorize memory systems for LM agents into two types: semantic and episodic. Semantic memory contains facts about the environment, and episodic memory stores past actions. In this work, we consider two baselines, Reflexion and Agent Workflow Memory (AWM), as exemplars of manipulating semantic and episodic memory [Shinn et al., 2023, Wang et al., 2025b]. Reflexion instructs the language model to reflect on the previous trajectory and propose areas of improvement; we consider these high-level notes about the environment to be part of semantic memory. AWM instructs the model to generate a summary workflow of the trajectory, provided the trajectory is successful; we consider this to be episodic memory. Building on this work, the current paper develops an improved mechanism for constructing and updating episodic memories.
@@
One reason why off-policy RL algorithms can be more efficient than on-policy ones is that they can store and learn from informative trajectories that are low probability under the current policy, or even trajectories that were never observed at all. Such experience replay techniques have proven especially valuable in situations with sparse rewards or limited data, as they extract maximal learning signal from a small number of positive examples [Schaul et al., 2016, Lu et al., 2023, Zhang et al., 2025]. In particular, hindsight experience replay (HER) further improves sample efficiency by relabeling past trajectories with alternative goals that were actually achieved during execution [Andrychowicz et al., 2017]. For instance, if an agent fails to reach a target location but successfully navigates to an intermediate point, HER treats this trajectory as a successful example for reaching that intermediate goal. Butt et al. [2024] apply HER to self-improve language models at writing code; our approach can be viewed as a generalization of HER in which not only goals, but arbitrary aspects of trajectories, can be edited in hindsight.
@@ 3.1 ECHO: Experience Consolidation via Hindsight Optimization
In the update rule, for each entity, we compare its newly generated descriptor to the descriptor’s predecessor and save the shorter workflow. Our motivation here is related to Kolmogorov complexity, or minimum description length (see Grünwald [2007] for an overview); we want the replay buffer to contain the shortest possible description for achieving the goal.
@@ 3.2 Baselines
In Listings 2 and 3, we cast Reflexion [Shinn et al., 2023] and AWM [Wang et al., 2025b] into our hindsight-update framework as well. The important parts of these algorithms lie in the hindsight rule, and the update rule simply appends the hindsights edits to memory.
@@ 4 Results
To test our algorithm, we create stateful versions of two environments where exploration is crucial: XMiniGrid [Nikulin et al., 2023] and PeopleJoinQA [Jhamtani et al., 2025]. These corresponding versions, XMiniGrid-Stateful and PeopleJoinQA-Stateful, can be used as benchmarks for efficient online learning in language agents; we release these environments at https://github.com/michahu/echo. We ran all experiments using GPT-4o [Hurst et al., 2024]; see Appendix 4 for hyperparameters.
@@ 4.1 Evaluation
where $R_{t}$ is the reward achieved in episode $t$. We use this metric to compare agents’ sample efficiency, as sample-efficient agents will rapidly increase the cumulative average reward. To normalize for problem difficulty, we report rewards as improvements over a baseline ReAct agent (Yao et al., 2023). Thus, the best method will be the one that maximizes both the final average reward and the rate at which it improves upon the ReAct agent.
@@ 5.2 Connections between RL and Prompting
ECHO continues the connection identified in ReACT [Yao et al., 2023] and Reflexion [Shinn et al., 2023] between reinforcement learning techniques and prompting strategies for language model agents. Tratitional RL methods like hindsight experience replay rely on numerical rewards and states, whereas language models can operate over experiences that can be described and modified through natural language. This enables a more flexible form of experience replay where the agent actively edits and improves past experiences based on its linguistic and commonsense understanding of the task. ECHO also bridges experience-based learning and symbolic reasoning, leveraging language models as both world models and policy generators.
@@ 5.3 Limitations and Future Work
In this work, we primarily considered natural language representations of semantic and episodic information. Recent work shows that code-like representations can be even more effective [Wang et al., 2025a], so future work should explore how ECHO’s performance changes when outputting programmatic trajectory representations. Furthermore, our update heuristic of accepting shorter trajectories for the same task can likely be improved in favor of a more sophisticated that combines new and old information while maintaining the same bias towards compression. Future work could also explore augmenting ECHO with retrieval-based mechanisms that draw from a memory bank (Zheng et al. [2024], Moghe et al. [2024]), enabling more targeted reuse of relevant experiences.
@@ A.2 PeopleJoinQA-Stateful
For the decision policy, we used the prompts from Jhamtani et al. [2025], available here: https://github.com/microsoft/peoplejoin.

### v2

@@ 1 Introduction
While language models (LMs) have demonstrated remarkable generalization across tasks, their performance often degrades in unfamiliar or interactive environments, especially when learning from limited experience [ziems-etal-2024-large, kwa2025measuringaiabilitycomplete, liang2023holistic]. In such settings, sample efficiency becomes critical, particularly when interactions are costly (e.g., with humans or physical systems). For example, a conversational assistant deployed for the first time in a new organization likely does not know where to look for specific pieces of information, or the best means of communicating with specific people. Thus, creating agents that can learn and adapt to their environments over time is of critical importance in improving their everyday usability.
Here we study the problem of building efficient mechanisms for online learning in LM agents. We consider the setting where an LM agent receives queries one at a time in a streaming fashion. Existing LM agent frameworks typically approach this setting through reflection [shinn2023reflexion, expel], memory [wang2025agent], or experience replay mechanisms [zheng2024synapse], which allow agents to revisit past episodes and improve over time. However, these methods primarily focus on storing or synthesizing experiences, and thus fail to fully exploit the LM’s ability to reason about counterfactuals—what could have led to success in past failures. This gap suggests an opportunity to design LM agents that actively rewrite and optimize their past experiences, converting failures into synthetic successes that improve future decision-making.
In this work, we introduce ECHO (Experience Consolidation via Hindsight Optimization), a framework that adapts hindsight experience replay (HER) to LM agents, enabling them to generate and learn from counterfactual trajectories for more sample-efficient learning. Our approach builds on hindsight experience replay methods from the RL literature [andrychowicz2017hindsight, HER]. HER learns a goal-conditioned policy; during training, each attempt to reach a goal state $s$ that fails in an end state $s^{\prime}$ is interpreted as a successful trajectory for reaching $s^{\prime}$. For example, a trajectory in which an LM fails to slice an apple by attempting to grab a knife, dropping it, then grasping an apple may still be interpreted as a demonstration of a successful grasp. But HER and related methods are relatively limited in the set of trajectory modifications they can make—relabeling trajectories with goals, but not altering the structure demonstrated trajectories themselves.
@@
We test ECHO and various state-of-the-art agent architectures in stateful variants of a 2D GridWorld task (XMiniGrid, nikulin2023xlandminigrid) and a question-answering task with multiple agents and tool calling (PeopleJoinQA, jhamtani-etal-2025-llm). These environments require exploration to successfully solve all queries of the task. In XMiniGrid, the agent must explore to find various objects in different rooms, and in PeopleJoin, the agent has imprecise or incomplete information about which teammates have the information it needs to answer a question. We make these environments stateful by allowing agents to persist insights via a scratchpad memory. Next, we reset the environment to the same initial position or state and vary the queries or tasks posed to the agent. The agent can then infer information about the environment over time and improve its performance and efficiency.
On XMiniGrid-Stateful, ECHO outperforms the baseline reason-then-act (ReAct) LM agent [yao2023react] by 80% in average reward and the next best baseline by 42%. On PeopleJoinQA-Stateful, ECHO still outperforms the standard ReAct agent on both accuracy and efficiency, while being slightly worse than the best baseline in accuracy by 4.6% and tied in efficiency. We conclude that ECHO is a promising technique for improving the sample efficiency of language agents, especially in environments where rewards are sparse and the baseline language agent performs poorly.
@@ 2 Related Work
Language model (LM) agents are systems that use large language models as reasoning engines to interact with environments, make decisions, and execute actions over time [yao2023react, schick2023toolformer]. These agents typically operate through a perception–action loop, where they observe their environment, reason about the current state, and generate actions [wang2023survey]. Recent work has demonstrated LM agents’ capabilities across diverse domains, from web navigation and tool use to multi-agent collaboration and code generation [liu2024agentbench, chatdev, jhamtani-etal-2025-llm]. However, a key limitation of current LM agents is their reliance on static knowledge encoded during pre-training, making them less effective when deployed in novel environments that require exploration and adaptation [zhou2024webarena]. This motivates the need for agents that can accumulate experience and improve their performance through interaction with their environment over time.
@@
Following sumers2024cognitive, we categorize memory systems for LM agents into two types: semantic and episodic. Semantic memory contains facts about the environment, and episodic memory stores past actions. In this work, we consider two baselines, Reflexion and Agent Workflow Memory (AWM), as exemplars of manipulating semantic and episodic memory [shinn2023reflexion, wang2025agent]. Reflexion instructs the language model to reflect on the previous trajectory and propose areas of improvement; we consider these high-level notes about the environment to be part of semantic memory. AWM instructs the model to generate a summary workflow of the trajectory, provided the trajectory is successful; we consider this to be episodic memory. Building on this work, the current paper develops an improved mechanism for constructing and updating episodic memories.
@@
One reason why off-policy RL algorithms can be more efficient than on-policy ones is that they can store and learn from informative trajectories that are low probability under the current policy, or even trajectories that were never observed at all. Such experience replay techniques have proven especially valuable in situations with sparse rewards or limited data, as they extract maximal learning signal from a small number of positive examples [PER, lu2023synthetic, zhang2025rlep]. In particular, hindsight experience replay (HER) further improves sample efficiency by relabeling past trajectories with alternative goals that were actually achieved during execution [andrychowicz2017hindsight]. For instance, if an agent fails to reach a target location but successfully navigates to an intermediate point, HER treats this trajectory as a successful example for reaching that intermediate goal. codeit apply HER to self-improve language models at writing code; our approach can be viewed as a generalization of HER in which not only goals, but arbitrary aspects of trajectories, can be edited in hindsight. bagel also take inspiration from HER to induce natural language descriptions of agent trajectories; we go one step further and induce synthetic trajectories themselves.
@@ 3.1 ECHO: Experience Consolidation via Hindsight Optimization
In the update rule, for each entity, we compare its newly generated descriptor to the descriptor’s predecessor and save the shorter workflow. Our motivation here is related to Kolmogorov complexity, or minimum description length (see grunwald2007minimum for an overview); we want the replay buffer to contain the shortest possible description for achieving the goal.
@@ 3.2 Baselines
In Listings 2 and 3, we cast Reflexion [shinn2023reflexion] and AWM [wang2025agent] into our hindsight-update framework as well. The important parts of these algorithms lie in the hindsight rule, and the update rule simply appends the hindsights edits to memory.
@@ 4 Results
To test our algorithm, we create stateful versions of two environments where exploration is crucial: XMiniGrid [nikulin2023xlandminigrid] and PeopleJoinQA [jhamtani-etal-2025-llm]. These corresponding versions, XMiniGrid-Stateful and PeopleJoinQA-Stateful, can be used as benchmarks for efficient online learning in language agents; we release these environments at https://github.com/michahu/echo. We ran all experiments using GPT-4o [hurst2024gpt]; see Appendix 4 for hyperparameters.
@@ 4.1 Evaluation
where $R_{t}$ is the reward achieved in episode $t$. We use this metric to compare agents’ sample efficiency, as sample-efficient agents will rapidly increase the cumulative average reward. To normalize for problem difficulty, we report rewards as improvements over a baseline ReAct agent (yao2023react). Thus, the best method will be the one that maximizes both the final average reward and the rate at which it improves upon the ReAct agent.
@@ 5.2 Connections between RL and Prompting
ECHO continues the connection identified in ReACT [yao2023react] and Reflexion [shinn2023reflexion] between reinforcement learning techniques and prompting strategies for language model agents. Tratitional RL methods like hindsight experience replay rely on numerical rewards and states, whereas language models can operate over experiences that can be described and modified through natural language. This enables a more flexible form of experience replay where the agent actively edits and improves past experiences based on its linguistic and commonsense understanding of the task. ECHO also bridges experience-based learning and symbolic reasoning, leveraging language models as both world models and policy generators.
@@ 5.3 Limitations and Future Work
In this work, we primarily considered natural language representations of semantic and episodic information. Recent work shows that code-like representations can be even more effective [wang2025asi], so future work should explore how ECHO’s performance changes when outputting programmatic trajectory representations. Furthermore, our update heuristic of accepting shorter trajectories for the same task can likely be improved in favor of a more sophisticated that combines new and old information while maintaining the same bias towards compression. Future work could also explore augmenting ECHO with retrieval-based mechanisms that draw from a memory bank (zheng2024synapse, moghe2024interpreting), enabling more targeted reuse of relevant experiences.
@@ A.2 PeopleJoinQA-Stateful
For the decision policy, we used the prompts from jhamtani-etal-2025-llm, available here: https://github.com/microsoft/peoplejoin.